# Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification

### v2

@@ 4.7 Effectiveness of the Noisy Synthetic Data Filter
Figure 8: Carotid stenosis and heart disease diagnostic accuracy (%) comparison using two classifiers on our filtering strategies under Real-finetune and Joint-train paradigms. The accuracy gains over the Baseline are depicted.
@@
Generative Training Paradigm Augmentation Method Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
/ None 75.51 0.718 59.28 0.532 80.55 0.681
/ TraAug 79.59 $\uparrow$ 4.08 0.737 $\uparrow$ 0.019 60.72 $\uparrow$ 1.44 0.559 $\uparrow$ 0.027 83.49 $\uparrow$ 2.94 0.733 $\uparrow$ 0.052
Real-finetune TraAug + GenAug 85.03 $\uparrow$ 9.52 0.813 $\uparrow$ 0.095 65.41 $\uparrow$ 6.13 0.611 $\uparrow$ 0.079 87.16 $\uparrow$ 6.61 0.795 $\uparrow$ 0.114
/ None 78.91 0.735 60.36 0.552 80.92 0.688
/ TraAug 80.95 $\uparrow$ 2.04 0.751 $\uparrow$ 0.016 61.08 $\uparrow$ 0.72 0.562 $\uparrow$ 0.010 81.28 $\uparrow$ 0.36 0.696 $\uparrow$ 0.008
Joint-train TraAug + GenAug 84.35 $\uparrow$ 5.44 0.811 $\uparrow$ 0.076 62.88 $\uparrow$ 2.52 0.578 $\uparrow$ 0.026 85.32 $\uparrow$ 4.40 0.762 $\uparrow$ 0.074
4.8 Impacts of Traditional and Generative Augmentations
We tested the contributions of different data augmentations in three downstream diagnostic tasks. The traditional augmentations used in this experiment can be found in Table 1. We report the performance gains of the two variants with data augmentations compared to the augmentation-free setup. For experimental fairness, we kept the training datasets for the None and TraAug variants the same size as the method combining traditional and generative augmentations. As presented in Table 6, the diagnostic performance shows noticeable improvements using the traditional augmentation method. Further coupled with the generative augmentation, the results show more substantial improvements than TraAug. We also observe that in the case of using traditional augmentation, adding extra real training samples yields limited performance gains, even a drop for ACDC. These prove that generative augmentation can effectively supplement traditional data augmentation and provide intrinsically diverse data characteristics for enhancing deep classifier learning, which support the viewpoints of (Luo et al., 2024).
@@ 5 Conclusion and discussion
In this study, we present a new and general diffusion-based generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification by leveraging customized and diagnosis-reliable synthetic sequences. To the best of our knowledge, this is the first comprehensive study to investigate the impact of the controllable generative augmentation scheme on aiding medical sequence classification.
@@
To improve the quality of synthetic data for promoting downstream classification, we propose a multimodal conditions-guided medical sequence generator that ensures flexibly controllable synthesis across semantic, sequential, and data distribution aspects. Moreover, we propose a highly effective noisy synthetic data filter to better learn from synthetic data, including adaptively filtering diagnosis-inhibitive synthetic sequences at class semantics and sequential levels. This can better connect the synthesis task and the downstream one, thus further enhancing our sequence classification performance. Extensive experiments on 3 datasets across multiple organs and modalities, with 11 popular networks trained on 3 paradigms, show the effectiveness and generality of Ctrl-GenAug. Furthermore, the empirical analysis validates that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness. We believe that Ctrl-GenAug can serve as a robust and practical data augmentation tool for various clinical scenarios.
@@
We then discuss the limitations of Ctrl-GenAug. First, the denoising diffusion implicit model (DDIM) sampling process is time-consuming due to the large inference steps. This limits the clinical practicality of Ctrl-GenAug for fast and large-scale production. Second, in out-domain conditions, the proposed framework requires target domain data during generator training, which can be tough to obtain in advance in clinical practice. In future work, we will adopt fast-sampling strategies (e.g., AMED-Solver (Zhou et al., 2024b)) to better balance the trade-off between sampling time and sample quality. Besides, we will incorporate test-time adaptation methods (Huang et al., 2023b; Luo et al., 2023; Huang et al., 2023c; Zhang et al., 2024b) to provide a more user-friendly tool for out-domain applications in clinical settings. Last, we will attempt to further extend the proposed method to segmentation and detection tasks (Huang et al., 2025; Tao et al., 2025).
@@
(Shang et al., 2023) disease class,
@@ 2 Fundus image
(Farooq et al., 2024) text classification CNN-based,
@@ 2 Dermoscopy image
(Zhang et al., 2024c) text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
(Sagers et al., 2023) text classification CNN-based 1 Dermoscopy image
(Khosravi et al., 2024) attribute, 14 pathology labels classification CNN-based 1 X-ray image
(Yu et al., 2024a) pathology labels,
@@
(Ktena et al., 2024) disease class, attribute classification CNN-based 1 Histopathology,
@@
(Luo et al., 2024) disease class classification CNN-based 10 Fundus,
@@
(Li et al., 2024a) multi-scale latent
@@
(Gong et al., 2024) mask segmentation CNN-based 1 MRI volume
(Peng et al., 2024) mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
(Kebaili et al., 2024) tumor attribute, mask segmentation CNN-based 1 MRI volume
(Yu et al., 2024b) mask segmentation CNN-based 1 CT volume
(Chen et al., 2024) mask segmentation CNN-based,
@@ 3 CT volume
(Dorjsembe et al., 2024a) mask segmentation CNN-based 1 MRI volume
(Huang et al., 2024a) mask segmentation CNN-based 2 OCT volume

### v3

@@ 4.8.3 Effectiveness of the Noisy Synthetic Data Filter
Table 8: Ablation study for performance-oriented diffusion settings in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018) datasets. I3D (Carreira and Zisserman, 2017) trained on Real-finetune paradigm was used for downstream diagnosis. The blue-highlighted block represents our default configuration.
Dataset Diffusion Settings Synthesis Diagnosis
@@
TUSC Cosine 12.5 5.19 92.79% 95.30% 63.24% 0.588
Cosine 7.5 3.90 92.57% 95.76% 64.69% 0.600
Scaled Linear 12.5 4.57 92.16% 95.45% 64.33% 0.597
Scaled Linear 7.5 3.69 92.14% 95.59% 65.41% 0.611
ACDC Cosine 12.5 8.42 73.95% 90.71% 85.32% 0.762
Cosine 7.5 8.63 73.88% 90.74% 86.05% 0.775
Scaled Linear 12.5 8.16 76.75% 92.56% 86.06% 0.774
Scaled Linear 7.5 8.18 77.08% 92.77% 87.16% 0.795
Table 9: Ablation study for efficiency-performance trade-off across diffusion settings in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018) datasets. I3D (Carreira and Zisserman, 2017) trained on Real-finetune paradigm was used for downstream diagnosis. All diffusion models are tested on a single NVIDIA GeForce RTX 3090 GPU. The blue-highlighted block represents our default configuration. DS, Dynamic Smoothness.
Dataset Diffusion Settings Synthesis Diagnosis
[…]
TUSC 3 200 13.26 86.29% 90.30% 53.5s 62.52% 0.575
3 300 12.33 83.42% 88.19% 87.4s 62.88% 0.580
4 100 3.70 92.02% 95.53% 39.8s 63.97% 0.592
4 200 3.69 92.14% 95.59% 68.8s 65.41% 0.611
4 300 2.69 91.15% 95.23% 100.3s 65.77% 0.622
ACDC 3 200 10.81 75.15% 88.73% 51.3s 83.85% 0.741
3 300 10.17 73.52% 88.41% 79.9s 83.85% 0.739
4 100 8.20 76.86% 92.74% 33.7s 85.69% 0.771
4 200 8.18 77.08% 92.77% 70.7s 87.16% 0.795
4 300 7.55 74.59% 91.90% 103.8s 88.06% 0.806
4.8.4 Impact of Diffusion Factors in Sequence Generator
To systematically analyze how the key diffusion factors affect the quality of the synthetic databases, we conducted two ablation studies that examine performance-oriented diffusion settings, including the noise schedule and guidance scale (Table 8), as well as efficiency-performance trade-off analyses under the settings of different UNet depths and sampling steps (Table 9).
As shown in Table 8, the scaled linear noise schedule consistently yields stronger downstream augmentation benefits and also achieves lower FVD than the cosine variant across both datasets. We believe that this advantage stems from its smoother and more stable noise evolution during training, which allows the model to more effectively learn and preserve anatomical details throughout the diffusion process. Moreover, the results show that using a high guidance scale of 12.5 leads to a noticeable degradation in diagnostic performance compared with the moderate setting of 7.5, with accuracy decreasing from 65.41% to 64.33% and AUROC dropping from 0.611 to 0.597 under the scaled linear schedule on the TUSC dataset. This trend suggests that excessive guidance may constrain the sampling trajectory, contributing to in-distribution (i.e., overly close to the real training distribution) synthetic samples, which in turn limit their usefulness for downstream model optimization. In contrast, a moderate guidance scale strikes a more favorable balance between semantic alignment and diversity, enabling the synthesizer to generate more diagnosis-promotive samples.
The depth of the UNet denoiser and the number of sampling steps both exert dual impacts on overall performance and synthesis efficiency. Table 9 demonstrates that increasing the UNet depth consistently improves generation fidelity and downstream classification, since deeper architectures can capture richer anatomical patterns and finer dynamic cues during training. However, these performance gains come with additional computational cost. A similar trend is observed for sampling steps, where additional iterations yield more diagnosis-reliable sequences but also require longer sampling time. Notably, configurations with insufficient feature representation capacity, such as a UNet depth of three, benefit minimally from increasing the sampling steps, indicating that additional steps cannot compensate for limited feature extraction ability. We further observe that the sequential consistency and smoothness of our synthetic samples are substantially more sensitive to the UNet depth than to other diffusion factors, as it directly determines the network’s capacity to model and propagate temporal/stereoscopic dependencies. Overall, our default configuration achieves a balanced compromise between generation quality and inference efficiency, providing a practical choice for large-scale customized medical sequence synthesis and generative augmentation.
Table 10: Sensitivity analysis on four representative classifier backbones used in the class semantics misalignment filtering across the Carotid, ACDC (Bernard et al., 2018), and MosMedData (Vladzymyrskyy et al., 2020) datasets. I3D (Carreira and Zisserman, 2017) trained on Real-finetune and Joint-train paradigms was adopted for downstream diagnosis. The accuracy and AUROC improvements over the Baseline are reported, with the best gains underlined.
Training Paradigm Semantic Filtering Classifier Carotid ACDC MosMedData
$\DeltaAccuracy Δ\DeltaAUROC$ $\DeltaAccuracy Δ\DeltaAUROC$ $\DeltaAccuracy Δ\DeltaAUROC$
Real-finetune CSN $\uparrow$ 4.08% $\uparrow$ 0.064 $\uparrow$ 2.56% $\uparrow$ 0.041 $\uparrow$ 4.76% $\uparrow$ 0.027
VideoSwin $\uparrow$ 4.08% $\uparrow$ 0.066 $\uparrow$ 1.83% $\uparrow$ 0.035 $\uparrow$ 5.66% $\uparrow$ 0.035
CSG-3DCT $\uparrow$ 5.44% $\uparrow$ 0.056 $\uparrow$ 2.57% $\uparrow$ 0.043 $\uparrow$ 4.25% $\uparrow$ 0.018
I3D $\uparrow$ 5.44% $\uparrow$ 0.076 $\uparrow$ 3.67% $\uparrow$ 0.062 $\uparrow$ 5.41% $\uparrow$ 0.043
Joint-train CSN $\uparrow$ 3.40% $\uparrow$ 0.042 $\uparrow$ 1.10% $\uparrow$ 0.021 $\uparrow$ 2.32% $\uparrow$ 0.007
VideoSwin $\uparrow$ 4.76% $\uparrow$ 0.045 $\uparrow$ 1.47% $\uparrow$ 0.027 $\uparrow$ 2.58% $\uparrow$ 0.005
CSG-3DCT $\uparrow$ 4.08% $\uparrow$ 0.071 $\uparrow$ 1.83% $\uparrow$ 0.028 $\uparrow$ 2.45% $\uparrow$ 0.006
I3D $\uparrow$ 4.76% $\uparrow$ 0.074 $\uparrow$ 1.83% $\uparrow$ 0.029 $\uparrow$ 2.32% $\uparrow$ 0.019
4.8.5 Analysis of Classifier Backbone Choice in the Semantic Filter
To investigate whether the observed diagnostic performance gains depend on the choice of classifier backbone used in the semantics misalignment filtering stage (refer to Sec. 6), we performed sensitivity analysis using four backbones across three datasets (Table 10). Specifically, we fixed the downstream classifier to I3D and selected four representative backbones in the semantic filtering stage. It can be observed that, under both training paradigms on three datasets, the diagnostic performance consistently improves with all backbone selections, and their gains exhibit only marginal variations (with deviations in accuracy gains of at most 1.84% relative to our default setting). Overall, these results demonstrate that the performance benefits are not tied to a particular semantic filtering backbone, affirming the robustness and general applicability of our method.
Figure 9: Diagnostic accuracy (%) comparison using I3D (Carreira and Zisserman, 2017) and FTC (Ahmadi et al., 2023) on three datasets under two annotation settings, both trained under the Real-finetune paradigm. Left: generator training with varying proportions of class-labeled real data. Right: downstream classifiers trained with different fractions of real labeled data. The gray dashed line indicates the Baseline result. Our Ctrl-GenAug can substantially reduce the overall annotation burden across both the generator training (Left) and downstream diagnosis stages (Right).
4.8.6 Analysis of Annotation Workload Reduction
To investigate the advantages of Ctrl-GenAug in reducing annotation burden, we conducted two dedicated ablation studies, including 1) examining the synthesizer’s dependency on real annotations during training (Figure 9, left), and 2) assessing whether the trained generator can reduce labeling demand in downstream applications (Figure 9, right). For the first study, we trained the generator using 25%, 50%, 75%, and 100% of the class-labeled real data, and subsequently used the corresponding synthetic datasets to compare their impact on downstream diagnostic accuracy under identical classifier settings. For the second study, we first obtained a synthetic dataset using the generator trained on all real labeled data, and then trained downstream classifiers with 25%, 50%, 75%, and 100% of the class-labeled real data under the Real-finetune paradigm to compare the resulting diagnostic accuracy. We also report the performance of classifiers trained exclusively on synthetic data (i.e., using 0% real annotated samples).
[…]
@@ 5 Conclusion and discussion
In this study, we present a new and general diffusion-based generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification by leveraging customized and diagnosis-reliable synthetic sequences. To the best of our knowledge, this is the first comprehensive empirical study on controllable generative augmentation for medical sequence classification.
@@
To improve the quality of synthetic data for promoting downstream classification, we propose a multimodal conditions-guided medical sequence generator that ensures flexibly controllable synthesis across semantic, sequential, and data distribution aspects. Moreover, we propose a highly effective noisy synthetic data filter to better learn from synthetic data, including adaptively filtering diagnosis-inhibitive synthetic sequences at class semantics and sequential levels. This can better connect the synthesis task and the downstream one, thus further enhancing our sequence classification performance. Extensive experiments on 5 medical datasets spanning 4 distinct modalities, including comparisons against 15 augmentation strategies and evaluations using 11 classifiers trained on 3 paradigms, comprehensively validate the robustness and generality of Ctrl-GenAug. Furthermore, the empirical analysis validates that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness. We believe that Ctrl-GenAug can serve as a powerful and practical data augmentation tool for various clinical scenarios.
Even though our experimental results demonstrate that synthetic data can effectively enhance medical AI models, we recognize that clinical adoption ultimately depends on rigorous validation showing that generated sequences are diagnosis-reliable and interpretable to radiologists. Expert evaluation provides an intuitive starting point, where radiologists assess fidelity, sequential coherence, and adherence to conditioning cues (Duan et al., 2025). Recording clear provenance for each synthetic sequence (such as its associated disease category and image priors) further enhances transparency and interpretability, which is particularly important for conditional generative methods. However, human review alone cannot scale to large synthetic databases, underscoring the need for automated and standardized quality-assessment pipelines. Building on prior insights (Huang et al., 2024c), we believe that a hierarchical and multidimensional evaluation suite tailored to medical sequence generation, complemented by small-scale user studies to ensure alignment between automated metrics and radiologists’ judgments, is urgently needed in the future to guarantee data quality and establish clinical credibility. This represents an interesting avenue for future exploration.
It is also worth mentioning that our proposed noisy synthetic data filter contributes an initial effort toward assessing the reliability of synthetic samples before being used in downstream diagnosis, and it can be seamlessly integrated into future large-scale quality-assessment pipelines. Ultimately, we suggest that synthetic data should complement rather than replace real clinical data, and its practical use relies on robust quality metrics, rigorous validation protocols, and appropriate regulatory oversight to ensure trustworthy and effective integration into medical AI workflows.
@@
We further discuss the limitations of Ctrl-GenAug. First, the denoising diffusion implicit model (DDIM) sampling process is time-consuming due to the large inference steps. This limits the clinical practicality of Ctrl-GenAug for fast and large-scale production. Second, in out-domain conditions, the proposed framework requires target domain data during generator training, which can be tough to obtain in advance in clinical practice. In future work, we will adopt fast-sampling strategies (e.g., AMED-Solver (Zhou et al., 2024b)) to better balance the trade-off between sampling time and sample quality. Besides, we will incorporate test-time adaptation methods (Huang et al., 2023b; Luo et al., 2023; Huang et al., 2023c; Zhang et al., 2024b) to provide a more user-friendly tool for out-domain applications in clinical settings. Last, we will further extend the proposed method to segmentation and detection tasks (Huang et al., 2025; Tao et al., 2025) and work toward establishing a comprehensive benchmark suite tailored for medical sequence generative models.
Data Availability Statement. Our data includes four publicly available datasets: TUSC (Stanford-AIMI, 2021), ACDC (Bernard et al., 2018), MosMedData (Vladzymyrskyy et al., 2020), and MRNet (Bien et al., 2018), as well as a private carotid dataset, which is available from the corresponding author upon reasonable request.
Acknowledgments. This work was supported by the grant from National Natural Science Foundation of China (Nos. 12326619, 62471313, 62471305), Science and Technology Planning Project of Guangdong Province (Nos. 2023A0505020002, 2024A1515030143), Frontier Technology Development Program of Jiangsu Province (No. BF2024078), Guangdong Basic and Applied Basic Research Foundation (Nos. 2025A1515011448, 2025A1515011821), Shenzhen Science and Technology Program (No. JCYJ20240813143302004), Shenzhen Natural Science Foundation (No. JCYJ20250604181708011), the Royal Academy of Engineering, United Kingdom under the RAEng Chair in Emerging Technologies (INSILEXCiET1919/19), the ERC Advanced Grant UKRI Frontier Research Guarantee (INSILICO EP/Y030494/1), the UK Centre of Excellence on in-silico Regulatory Scienceand Innovation (UK CEiRSI) (10139527), the National Institute for Health and CareResearch (NIHR) Manchester Biomedical Research Centre (BRC) (NIHR203308), the BHF Manchester Centre of Research Excellence (RE/24/130017), and the CRUKRadNet Manchester (C1994/A28701).
@@
Figure S2: Additional synthetic results on MosMedData (Vladzymyrskyy et al., 2020) (a-d) and MRNet (Bien et al., 2018) (e-h). We used the multimodal conditions joint training strategy (Huang et al., 2023a) during training and flexibly dropped several during sampling.
@@
Shang et al. (2023) disease class,
@@ 2 Fundus image
Farooq et al. (2024) text classification CNN-based,
@@ 2 Dermoscopy image
Zhang et al. (2024c) text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
Sagers et al. (2023) text classification CNN-based 1 Dermoscopy image
Khosravi et al. (2024) attribute, 14 pathology labels classification CNN-based 1 X-ray image
Yu et al. (2024a) pathology labels,
@@
Ktena et al. (2024) disease class, attribute classification CNN-based 1 Histopathology,
@@
Luo et al. (2024) disease class classification CNN-based 10 Fundus,
@@
Li et al. (2024a) multi-scale latent
@@
Gong et al. (2024) mask segmentation CNN-based 1 MRI volume
Peng et al. (2024) mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
Kebaili et al. (2024) tumor attribute, mask segmentation CNN-based 1 MRI volume
Yu et al. (2024b) mask segmentation CNN-based 1 CT volume
Chen et al. (2024) mask segmentation CNN-based,
@@ 3 CT volume
Dorjsembe et al. (2024a) mask segmentation CNN-based 1 MRI volume
Huang et al. (2024a) mask segmentation CNN-based 2 OCT volume